If a=sin(pi/4), b=cos(pi/4) and c=-cosec(pi/4) then the value of a³+b³+c³ is

Grade:10

2 Answers

046 Nirbhay Gupta
2 years ago
a = sin 45 = 1/√2
b = cos 45 = 1/√2
c = -cosec 45 = -√2
a3+b3+c3 = (1/√2)3 + (1/√2)3 + (-√2)3 
                = (1/2√2) + (1/2√2) – (2√2)
                = -3/√2
Mukul
one year ago
 sin 45 = 1/√2,  cos 45 = 1/√2,  -cosec 45 = -√2
a^3+b^3+c^3 = (1/√2)^3 + (1/√2)^3 + (-√2)^3 
                         = (1/2√2) + (1/2√2) – (2√2)
                         = -3/√2